# Tree Proof-of-Position Algorithms

### v1

@@ XI Conclusion
In summary, we provide an algorithm for agents to prove their position in a decentralised, collaborative and privacy preserving manner. Our algorithm is robust to powerful adversaries, and we make no assumption of honest behaviour in the system. We provide a simple mathematical model to aid users to select the most suitable operating conditions of T-PoP, provided the security and reliability guarantees they wish to achieve, and the expected agent density in their system. We validate the model through the use of extensive Monte-Carlo agent-based simulations, and find these are in complete agreement with our model. Our code base is open-source, and the results along with the mathematical model and the agent-based simulator can be found in our GitHub Repository. For future work, we consider implementing this protocol in useful applications.

### v2

@@ XI Conclusion
In this paper, we provide an algorithm for agents to prove their position in a decentralised, collaborative and privacy preserving manner. The analysis focuses on the robustness of the algorithm in an adversarial environment, showing that TPoP is robust to adversarial attacks. We provide a simple mathematical model to aid users to select the most suitable operating conditions of T-PoP, provided the security and reliability guarantees they wish to achieve, and the expected agent density in their system. The mathematical model is validated through the use of extensive Monte-Carlo agent-based simulations. Our code base is open-source, and the results along with the mathematical model and the agent-based simulator can be found in our GitHub Repository.